OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have

OpenStudy (anonymous):

Number of quarters = q 25q + 10(46 - q) = 700 5q + 2(46 - q) = 140 5q + 92 - 2q = 140 3q = 48 q = 16 No. of dimes: = 46 - 16 = 30 Answer: 16 quarters & 30 dimes